OpenStudy (anonymous):

A bag contains 6 cherry candies, 3 orange candies, and 2 lemon candies. You reach in and take 3 pieces of candy at random. What is the probability that you have at least 1 lemon candy? Round to the nearest ten-thousandth when necessary.

OpenStudy (anonymous):

at least one means not none

OpenStudy (anonymous):

you can compute the probability you get no lemon candies, then subtract that from one or you can compute the probability you get one lemon, and the probability you get two lemon, and add them up which would you like to do?

OpenStudy (anonymous):

the second option

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

at first thought it was 2C1/11C3 but it didn't work out.

OpenStudy (anonymous):

number of ways to choose 3 out of 11 is \(\binom{11}{3}\) which you may have seen written as \(_{11}C_3\) and is computed via \[\frac{11\times 10\times 9}{3\times 2}=165\] and that will be your denominator

OpenStudy (anonymous):

oh i see, ok since you have a good idea, lets do the easy way the probability you get no lemon candies is \[\frac{_9C_3}{_{11}C_3}\]

OpenStudy (anonymous):

find that number, subtract it from 1, and you are done

OpenStudy (anonymous):

these r the answer choices I have

OpenStudy (anonymous):

0.1515 0.3362 0.4909 0.5758

OpenStudy (anonymous):

ok I see

OpenStudy (anonymous):

if you want the probability you have exactly one, you have to take in to account that it is one lemon, and two not lemon, so it would be \[\frac{_9C_2\times 2}{_{11}C_3}\]

OpenStudy (anonymous):

and then you can compute the probabilty you have exactly 2, which is \[\frac{_9C_1}{_{11}C_3}\] or just \(\frac{9}{165}\) and then you have to add to the first number, but the other method is easier

OpenStudy (anonymous):

thank you very much for ur help

OpenStudy (anonymous):

Another approach, based on P(0) idea. first pick, 2/11 lemon 1-2/11=0.82 probability of not lemon second pick 2/10 lemon 1-2/10 = 0.80 third ick, 2/9 lemon 1-2/9 = 0.78 Multiply these probs get P(0) in three picks = 0.509 Thus prob at least one is 0.491

OpenStudy (anonymous):

thank you